# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v4

@@ Abstract
Recent advancements in real-time neural rendering using point-based techniques have paved the way for the widespread adoption of 3D representations. However, foundational approaches like 3D Gaussian Splatting come with a substantial storage overhead caused by growing the SfM points to millions, often demanding gigabyte-level disk space for a single unbounded scene, posing significant scalability challenges and hindering the splatting efficiency. To address this challenge, we introduce LightGaussian, a novel method designed to transform 3D Gaussians into a more efficient and compact format. Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects. Additionally, LightGaussian employs distillation and pseudo-view augmentation to distill spherical harmonics to a lower degree, allowing knowledge transfer to more compact representations while maintaining scene appearance. Furthermore, we propose a hybrid scheme, VecTree Quantization, to quantize all attributes, resulting in lower bitwidth representations with minimal accuracy losses. In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 139 to 215, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets. Project website: https://lightgaussian.github.io/
Figure 1: Compressibility and Rendering Speed. We present LightGaussian to transform 3D Gaussians into a more compact representation. LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left). Consequently, it reduces the average storage from 727MB to 42MB and improves the FPS from 139 to 215.
11footnotetext: Equal Contribution.
@@ 1 Introduction
Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images. NVS holds importance as it can be used for a wide range of real-world applications, including virtual reality [12], augmented reality [71], digital twin [13], and autonomous driving [63]. Neural Radiance Fields (NeRFs)[43, 4, 5] have demonstrated promising ability for photo-realistic 3D modeling and synthesis from multi-view images where 3D location and view directions are mapped to view-dependent color and volumetric density. The pixel intensity can be rendered using the volume rendering technique[14]. However, NeRF and its variants have struggled with rendering speeds for practical deployment in real-world scenarios. Subsequent efforts along this trend introduce voxel-based implicit representation [55, 62, 51, 67, 10], hash grid [45], computational parallelism [48], or neural light field representation [60] to speed up the rendering. However, they either require task-specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed. Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes. The authors utilize the idea of representing the scene as explicit 3D Gaussians with specific properties to model the scene; the 2D images are efficiently rendered using a technique named splatting[36]. The optimal balance between speed and quality indicates a potential trend to use 3D-GS as a new representation for generating numerous large-scale scenes in digital twins and autonomous driving. However, using point-based representations inevitably introduces significant storage costs, as each point and its attributes are stored independently. For instance, 3D-GS on a typical unbounded 360 scene [5] requires more than one gigabyte, which inhibits the scalability of 3D-GS (e.g., 1.4GB on scene Bicycle).
@@
In this paper, we counteract the heavy storage issue and deliver a compact representation while preserving the rendering quality. By examining the well-trained point-based representation, each scene is composed of millions of Gaussians, grown from sparse Structure from Motion (SfM) point cloud. Attributes are attached to Gaussians to model the scene’s geometry and appearance. However, the significant number of Gaussians, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes. To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis. Gaussians with low impact on visual quality will be identified and pruned, followed by short recovery steps to be applied. Spherical harmonics (SH) coefficients, as the majority of the data, are used for modeling the view-dependent color. Compressing it directly by shrinking the higher degree harms the reflectance effect, and we propose a general distillation step enhanced by pseudo-view augmentation to harmlessly transfer the knowledge into a compact level. VecTree Quantization step adaptively picks “just some amount” of distinct point attributes based on the global significance, further reducing the required bitwidth of the original format.
@@
In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig. 1, left), significantly reducing storage requirements from 727MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.013 in SSIM) on the Mip-NeRF 360 datasets (Fig. 1, right). LightGaussian further improves the rendering speed to a higher level (200+ FPS) on complex scenes containing detailed backgrounds, suggesting a viable solution for broadening the application scope.
@@ 2 Related Works
Figure 2: The overall pipeline of LightGaussian. 3D Gaussians are optimized from multi-view images and SfM points. LightGaussian first calculates the global significance for each Gaussian based on training observations; Gaussians with the least significance are pruned. A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format. VecTree quantization, consisting of lossy codebook quantization and lossless octree quantization, is further introduced for compressing the representation.
Efficient 3D Scene Representations for NVS
Neural radiance field (NeRF) [43] uses a multi-layer perceptron (MLP) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [66, 21, 39, 31, 55, 45, 10], caching and distillation [60, 22, 28, 48], or ray-based representations [2, 54] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [46] and cloud rendering applications [41]. Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques [72] for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5]. Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
Model Pruning and Vector Quantization
@@
3 Method
Overview
The overview of LightGaussian is illustrated in Fig. 2. The 3D-GS model is trained using multi-view images and is initially initialized from SfM point clouds. By expanding the sparse points to millions of Gaussians, the scene is well-represented. Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format. This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the modeled specular light, and VecTree Quantization to store Gaussians at a lower bit-width.
@@ 3.1 Background: 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) [32] is an explicit point-based 3D scene representation, utilizing Gaussians with various attributes to model the scene. When representing a complex real-world scene, 3D-GS is initialized from an SfM sparse point cloud, and Gaussian Densifications are applied to increase the Gaussian counts that are used for handling small-scale geometry insufficiently covered. Formally, each Gaussian is characterized by a covariance matrix $\boldsymbol{\Sigma}$ and a center point $\boldsymbol{X}$, which is referred to as the mean value of the Gaussian:
@@
When rendering 2D images from the 3D Gaussians, the technique of splatting [65, 36] is employed for the Gaussians within the camera planes. With a viewing transform denoted as $\boldsymbol{W}$ and the Jacobian of the affine approximation of the projective transformation represented by $\boldsymbol{J}$, the covariance matrix $\boldsymbol{\Sigma}^{\prime}$ in camera coordinates can be computed as follows:
$\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}.$ (2)
Specifically, for each pixel, the color and opacity of all the Gaussians are computed using the Gaussian’s representation Eq. 1. The blending of $N$ ordered points that overlap the pixel is given by the formula:
$\boldsymbol{C}=\sum_{i\in N}\boldsymbol{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients. In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
Figure 3: Zero-shot Opacity-based Pruning. A significant number of Gaussians exhibit small opacity values (top). Simply utilizing Gaussian opacity as an indicator for pruning the least important Gaussians results in the rendered image losing intricate details (bottom), with the PSNR dropping from 27.2 to 25.3. This has inspired us to find better criteria to measure global significance in terms of rendering quality. The accumulated PDF is equal to 1.
Global Significance Calculation
Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula. Inspired by Eq 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering. The initial significance score of each Gaussian can then be quantified based on its contribution to each pixel (a.k.a rays, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian. The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and opacity $\sigma_{j}$. The volume calculation equation is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$). We implement this in PyTorch as torch.prod(gaussians.scaling, dim=1) without considering the constant value $\frac{4}{3}\pi$. Overall, the global significant score can be summarized as:
@@
where $j$ indicates the Gaussian index, and $M$, $H$, and $W$ represent the number of training views, image height, and width, respectively. $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
@@
Here, the volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, and $\beta$ is introduced to provide additional flexibility.
Gaussian Co-adaptation
We rank all the Gaussians based on the computed global significance scores, and provide a quantitative basis for pruning Gaussians with the lowest scores. It is worth noting that an overly aggressive pruning ratio can be detrimental to performance. Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage. To mitigate this, we propose to utilize photometric loss on original training views to optimize (adapt) the remaining Gaussians, which jointly adjusts their attributes over 5k steps, without further densification, which can compensate for the loss incurred by pruning.

### v5

@@ Abstract
Recent advancements in real-time neural rendering using point-based techniques have paved the way for the widespread adoption of 3D representations. However, foundational approaches like 3D Gaussian Splatting come with a substantial storage overhead caused by growing the SfM points to millions, often demanding gigabyte-level disk space for a single unbounded scene, posing significant scalability challenges and hindering the splatting efficiency. To address this challenge, we introduce LightGaussian, a novel method designed to transform 3D Gaussians into a more efficient and compact format. Drawing inspiration from the concept of Network Pruning, LightGaussian identifies Gaussians that are insignificant in contributing to the scene reconstruction and adopts a pruning and recovery process, effectively reducing redundancy in Gaussian counts while preserving visual effects. Additionally, LightGaussian employs knowledge distillation and pseudo-view augmentation to transfer spherical harmonics efficients to a lower degree, allowing knowledge convert to more compact representations while maintaining scene appearance. Furthermore, we propose a Gaussian Vector Quantization based on Gaussian global significance, to quantize all redundant attributes, resulting in lower bitwidth representations with minimal accuracy losses. In summary, LightGaussian achieves an averaged compression rate over 15$\times$ while boosting the FPS from 119 to 209, enabling an efficient representation of complex scenes on Mip-NeRF 360, Tank & Temple datasets.
Figure 1: Compressibility and Rendering Speed. We present LightGaussian to transform 3D Gaussians into a more compact representation. LightGaussian effectively prunes redundant Gaussians while preserving visual fidelity (on the left). It reduces the average storage from 724MB to 42MB and improves the FPS from 119 to 209.
@@ 1 Introduction
Novel view synthesis (NVS) aims to generate photo-realistic images of a 3D scene from unobserved viewpoints, given a set of calibrated multi-view images. NVS holds importance as it can be used for a wide range of real-world applications, including virtual reality [12], augmented reality [71], digital twin [13], and autonomous driving [64]. Neural Radiance Fields (NeRFs)[45, 4, 5] have demonstrated promising ability for photo-realistic 3D modeling and synthesis from multi-view images where 3D location and view directions are mapped to view-dependent color and volumetric density. The pixel intensity can be rendered using the volume rendering technique[14]. However, NeRF and its variants have struggled with rendering speeds for practical deployment in real-world scenarios. Subsequent efforts along this trend introduce voxel-based implicit representation [57, 62, 54, 67, 10], hash grid [46], computational parallelism [51], or neural light field representation [61] to speed up the rendering. However, they either require specific design or face a trade-off between rendering quality and efficiency, making them difficult to generalize to large-scale scenarios with practical speed. Recent progress in point-based 3D Gaussian Splatting (3D-GS) [32] has brought photo-realistic rendering quality to the real-time level, even for complex scenes. The authors utilize the idea of representing the scene as explicit 3D Gaussians with specific properties to model the scene; the 2D images are efficiently rendered using a technique named splatting[36]. The optimal balance between speed and quality indicates a potential trend to use 3D-GS as a new representation for generating numerous large-scale scenes in digital twins and autonomous driving. However, using point-based representations inevitably introduces significant storage costs, as each point and its attributes are stored independently. For instance, 3D-GS on a typical unbounded 360 scene [5] requires more than one gigabyte, which inhibits the scalability of 3D-GS (e.g., 1.4GB on scene Bicycle).
@@
In this paper, we counteract the heavy storage issue and deliver a compact representation while preserving the rendering quality. By examining the well-trained point-based representation, each scene is composed of millions of Gaussians, grown from sparse Structure from Motion (SfM) point cloud. Attributes are attached to Gaussians to model the scene’s geometry and appearance. However, the significant number of Gaussians after its optimization, as well as the high-degree Spherical Harmonics (SH) coefficients used for modeling scene reflection, contribute to an over-parametrized representation when fitting to the scenes. To minimize the required Gaussian number, we propose proper criteria to measure the global significance of each 3D Gaussian, in the context of its contribution for view synthesis. Gaussians that have minimal impact on visual quality are identified and pruned. This is followed by streamlined recovery steps that allow for rapid optimization, named as Gaussian Pruning and Recovery. Spherical harmonics (SH) coefficients, constituting a significant portion of the data, are employed to model view-dependent color and to capture the interaction of light with surfaces from diverse angles. However, utilizing higher-degree SH coefficients presents a trade-off between accuracy and computational efficiency. While higher-degree coefficients offer more detailed lighting information, enhancing scene realism, they also escalate computational complexity and storage demands. To mitigate this, we propose a SH Distillation step complemented by pseudo-view augmentation. This approach is designed to efficiently condense the information from higher-degree coefficients into a more compact form, ensuring that quality is preserved without inflating computational requirements. Besides focusing on reducing redundancy in spatial (Gaussian count) and lighting (SHs) representations, we additionally introduce a Vector Quantization step. This process adaptively selects a specific quantity of distinct point attributes, determined by their global significance, which further reduces the required bitwidth from the original format.
@@
In summary, our proposed framework, LightGaussian, efficiently reduces the Gaussian count (e.g., from 1.49M to 575K, Fig. 1, left), significantly reducing storage requirements from 724MB to 42MB, with minimal rendering quality decrease ($\downarrow$ 0.005 in SSIM) on the Mip-NeRF 360 dataset (Fig. 1, right). LightGaussian further improves the rendering speed to a higher level (200+ FPS) on complex scenes containing detailed backgrounds, suggesting a viable solution for broadening the application scope.
@@ 2 Related Works
Efficient 3D Scene Representations for NVS.
Neural radiance field
(NeRF) [45] propose to use multi-layer perceptron (MLPs) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [68, 21, 41, 31, 57, 46, 10], caching and distillation [61, 22, 28, 51], or ray-based representations [2, 56] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [49] and cloud rendering applications [43]. Recently, Pulsar [37] has shown its efficient sphere rasterization, 3D Gaussian Splatting (3D-GS) [32] utilizes anisotropic Gaussians [72] and introduces a tile-based sorting for rendering, achieves real-time speed with comparable rendering quality to the MLP-based ones(e.g., Mip-NeRF 360 [5]). Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS. Concurrent works adopt region-based vector quantization [48], K-means based codebooks [47], or learned binary masks for each Gaussian, replacing the SHs with grid-based neural networks [39] within a modified training framework to reduce model size.
Model Pruning and Vector Quantization.
@@
3 Methods
Figure 2: The overall pipeline of LightGaussian. 3D Gaussians are optimized from multi-view images and SfM points. LightGaussian first calculates the global significance for each Gaussian on training data, where Gaussians with the least significance are pruned. A subsequent distillation with synthesized pseudo-views is introduced to transfer the SH into a compact format. Vector quantization, consisting of codebook initialization and assignment, is further introduced for reducing the model bandwidth.
Overview.
The overview of LightGaussian is illustrated in Fig. 2. The 3D-GS model is trained using multi-view images and is initialized from SfM point clouds. By expanding the sparse points to millions of Gaussians, the scene is well-represented. Then, the 3D-GS undergoes processing within our pipeline to transform it into a more compact format. This involves utilizing Gaussian Prune and Recovery to reduce the number of Gaussians, SH Distillation to remove redundant SHs while preserving the essential lighting information, and Vector Quantization to store Gaussians at a lower bit-width.
@@ 3.1 Background: 3D Gaussian Splatting
3D Gaussian Splatting (3D-GS) [32] is an explicit point-based 3D scene representation, utilizing Gaussians with various attributes to model the scene. When representing a complex real-world scene, 3D-GS is initialized from a sparse point cloud generated by SfM, and Gaussian Densification is applied to increase the Gaussian counts that are used for handling small-scale geometry insufficiently covered and over reconstruction. Formally, each Gaussian is characterized by a covariance matrix $\boldsymbol{\Sigma}$ and a center point $\boldsymbol{X}$, which is referred to as the mean value of the Gaussian:
@@
When rendering 2D images from the 3D Gaussians, the technique of splatting [66, 36] is employed for the Gaussians within the camera planes. With a viewing transform denoted as $\boldsymbol{W}$ and the Jacobian of the affine approximation of the projective transformation represented by $\boldsymbol{J}$, the covariance matrix $\boldsymbol{\Sigma}^{\prime}$ in camera coordinates can be computed as $\boldsymbol{\Sigma}^{\prime}=\boldsymbol{JW}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T}$. Specifically, for each pixel, the color and opacity of all the Gaussians are computed using the Gaussian’s representation Eq. 1. The blending of $N$ ordered points that overlap the pixel is given by the formula:
$\boldsymbol{C}=\sum_{i\in N}{c_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (2)
Here, $c_{i}$, $\alpha_{i}$ represents the view-dependent color and opacity, calculated from a 2D Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-3D Gaussian’s opacity. In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\text{SH}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminates less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for 3D Gaussian Splatting representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
Figure 3: Zero-shot Opacity-based Pruning. A significant number of Gaussians exhibit small opacity values (top). Simply utilizing Gaussian opacity as an indicator for pruning the least important Gaussians results in the rendered image losing intricate details (bottom), with the PSNR dropping from 27.2 to 25.3. This has inspired us to find better criteria to measure global significance in terms of rendering quality. The accumulated Probability Density Function(PDF) is equal to 1.
Global Significance Calculation.
Rather than merely depending on Gaussian opacity for assessing significance, we are inspired by Eq. 2. In this context, 3D Gaussians within the view frustum are evaluated and rasterized by projecting them onto a camera viewpoint for rendering. The initial significance score of each Gaussian (denoted as $\text{G}(\boldsymbol{X}_{j})$) can then be quantified based on its contribution to each pixel (ray, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$ whether they intersect or not. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian. The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and 3D Gaussian’s opacity $\sigma_{j}$, as they all contribute for the rendering formula. The volume calculation equation of the $j$-th Guassian is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$). Finally, the global significant score can be summarized as:
@@
where $j$ is the Gaussian index, $i$ means a pixel, $M$, $H$, and $W$ represents the number of training views, image height, and width, respectively. $\mathbbm{1}$ is the indicator function that determines whether a Gaussian intersects with a given ray.
@@
Here, the calculated Gaussian volume is firstly normalized by the 90% largest of all sorted Gaussians, clipping the range between 0 and 1, to avoid excessive floating Gaussians derived from vanilla 3D-GS. The $\beta$ is introduced to provide additional flexibility.
Gaussian Co-adaptation.
We rank all the Gaussians based on the computed global significance scores, which provides a quantitative basis for pruning Gaussians with lower scores. The remaining ones are optimized (adapted) to jointly adjust their attributes, without further densification, to compensate for the slight loss incurred by pruning. Note that, we utilize photometric loss on original training views with 5,000 iterations which is aligned with 3D-GS training.